solve the following equation (show work)


10/x = 5 (10 over x = 5)


1
Expert's answer
2021-05-07T10:07:29-0400

10/x =5

multiply both sides by denominator to remove fraction on LHS:

x*(10/x) = 5 *x


10=5x

divide through by 5 to get value of x on RHS:

(10) /5 = (5x) /5


2= x the value of x